# The SABRE South Technical Design Report Executive Summary

### v2

@@ 1 Introduction
Summarised in this document is the Technical Design Report for the SABRE (Sodium Iodide with Active Background REjection) South experiment. SABRE South is a NaI(Tl) dark matter direct detection experiment, to be situated underground in the Stawell Underground Physics Laboratory, and aims to test the results from the DAMA/LIBRA experiment. SABRE South’s three key sub-detectors are described, alongside other major components of the experiment design such as data acquisition, and shielding. The expected sensitivity of the detector to a DAMA/LIBRA-like signal is also discussed.
@@ 1.1 Scientific goals
A wide range of astrophysical evidence supports the existence of non-luminous matter known as dark matter (DM). Based on these observations, dark matter must have the following properties: it interacts gravitationally, but is electrically neutral and does not interact via the strong force; it is largely non-relativistic (“cold”); it is around five times more abundant than baryonic matter; it has a freezeout time that allows for the observed structural evolution of the universe; and, it is stable on cosmological timescales. Any observable signal from direct detection should have a distinct modulation signature due to the Earth’s revolution around the Sun as it moves through the galactic dark matter halo. This modulation is unrelated to the Earth’s seasons and is identical in both hemispheres. To be compatible with DM, the modulation signal should satisfy the following conditions: the modulation should have a period of one year; the modulation’s phase should produce a peak at the start of June; for GeV-scale dark matter particle(s), the modulation should be present in a low energy range (keV-scale); the dark matter particle interacts rarely with matter, expecting only single hits with the target material; the modulation amplitude is not larger than $\sim$30of the total dark matter signal (the exact amount varies depending on the particle interaction model).
@@ 1.2 Dark Matter searches in NaI(Tl) detectors
DAMA/LIBRA, like its predecessor DAMA/NaI, aims to observe the model-independent modulation of a dark matter interaction rate. The most recent iteration, DAMA/LIBRA-phase 2 (and phase 1), used 250 kg of NaI(Tl) over 8 years of data collection. Combining this with previous experiments, DAMA/LIBRA has a total exposure period of 2.86 tonne years. The interaction rate itself, integrated over a given energy window $k$, can be given by
@@
where $R_{0,k}$ is the constant rate, $S_{m,k}$ is the amplitude of the modulating component of the rate, $\omega$ is the angular frequency, and $t_{0}$ is the phase [2].
@@
The DAMA collaboration has consistently observed a modulating signal over their background rate of 0.8 cpd/kg/keV [2], compatible with a dark matter signal over the DAMA/NaI, DAMA/LIBRA phase 1 and DAMA/LIBRA phase 2 experiments in the 2-6 keV region, this has a modulation amplitude of $0.01014\pm 0.00074$ cpd/kg/keV with a phase and period of 142.4 $\pm$ 4.2 days and 0.99834 $\pm$ 0.00067 years respectively [3]. In the 1-6 keV recoil energy range (reported for DAMA/LIBRA-phase 2 only) this modulation occurs with an amplitude of $0.01058\pm 0.00090$ cpd/kg/keV, period of $0.99882\pm 0.00065$ years and phase of $144.5\pm 5.1$ days. Recent preliminary results indicate that this modulation persists below the 1 keV threshold [1].
@@
The signal observed by DAMA/LIBRA is in direct tension with a wide variety of direct-detection dark matter experiments using different target materials [4]. Recently, there have also been studies suggesting that DAMA/LIBRA’s background averaging technique (if performed over a time-dependent background) could have induced a modulating result [5]. However, after reproducing a best estimate of their background, we find that: (i) the reproduced background is too low for a modulation to be induced on the same scale as their reported signal, and (ii) the resulting induced modulation is opposite in phase to the modulation reported by DAMA/LIBRA [6]. Therefore, in recent years, other NaI(Tl)-based experiments have been developed to test this modulation using the same target as DAMA, such as COSINE [7] and ANAIS [8]. We list the observed modulation from each of the operating experiments in Table 1. Results in the 1-6 keV region for COSINE and ANAIS are known to be impacted by higher than expected background contributions that are unexplained [9, 8]; a potential cause of this is photomultiplier noise [10]. Additionally, the signal efficiencies of the event selection criteria drop steeply as 1 keV is approached. For example, between 1-2 keV, the ANAIS detection efficiency drops to as low as 20%, with the degree of background rejection not reported [8]. As a result, a comparison of results in the 2-6 keV region is likely more robust. To definitively test DAMALIBRA, SABRE must be sensitive to a modulation of 0.01 cpd / kg / keV, which requires a total background of no more than 1 cpd/kg/keV, and ideally better than the DAMA value of 0.8 cpd/kg/keV [2].
@@
Table 1: Annual modulation results reported by DAMA [3], COSINE [9], and ANAIS [11] (ANAIS also reports updated results from their 3-year data [8]). We define dru (differential rate units) as cpd/kg/keV.
Setup Mass (kg) Bkg. (dru) $S_{m}$ (1-6 keV) (dru) $S_{m}$ (2-6 keV) (dru)
DAMA/LIBRA phase 2 250 0.8 0.0105 $\pm$ 0.00090 0.0093 $\pm$ 0.00094
COSINE 61.3 2.7 0.0067 $\pm$ 0.0042 0.0051 $\pm$ 0.0047
@@ 2 The Stawell Underground Physics Laboratory
Figure 1: Left: location of SUPL. Right: muon flux at a number of underground laboratories as a function of water equivalent depth. The point for SUPL is based on simulated results using GEANT4 and MUTE [12].
The construction of SUPL was completed in 2023, with the first access for SABRE South installation activities in early 2024. Assembly and commissioning of the main detector components in SUPL is currently scheduled for 2025, after which data collection will begin. SUPL has helical drive access through mining tunnels, which has implications for how large detector components are transported and designed.
@@ 3 The SABRE South concept
To definitively test the DAMA/LIBRA anomaly, not only does the same target material need to be used, but an ultra low background is required in order to be sensitive to the 0.01 cpd/kg/keV reported modulation amplitude. With the main component of the total background being the radiogenic background in the crystal, we aim to bring the background of the SABRE South experiment to levels comparable with those of DAMA/LIBRA by ensuring high NaI(Tl) radiopurity. Furthermore, by employing a background veto system, in the form of both the liquid scintillator detector and muon detector, external background and remaining radioimpurities can be tagged and removed. The test crystals grown for the SABRE collaboration are competitive with similar NaI(Tl) in terms of their radiopurity [13, 14, 15].
@@ 3.1 Radiopurity of NaI(Tl) detectors
As indicated in Refs. [14, 13], and also shown by SABRE South simulations in Ref. [16], radiogenic backgrounds intrinsic to the crystals dominate the total background, meaning radiopure crystals are the key factor in achieving a low-background experiment. The purity levels of the DAMA, ANAIS, COSINE, and SABRE crystals are listed in Table 2, with the predicted SABRE South energy spectrum (with the veto efficiency applied) shown in Fig. 3. Some of the background contributions are suppressed by the active veto system, while others need to be reduced through purification processes applied during the growth of the crystals. The SABRE crystals already offer a large reduction in 40K and, with the exception of DAMA, have the lowest levels of 210Pb. Further crystal production consistent with these results would see SABRE becoming the lowest-background NaI(Tl) dark matter detector worldwide, and is, by design, the driving factor behind the 0.72 cpd/kg/keV total background reported in Ref. [16]. Further details on NaI (Tl) background modelling for SABRE South can be found in Ref. [16].
@@
SABRE (NaI-033) [15] $4.3$ $0.4$ $0.2$ $0.34$
Figure 3: The expected crystal background rate from intrinsic contamination at SABRE South for 50 kg of NaI(Tl) crystals [16].
@@ 3.2 SABRE South detector
The SABRE South experiment is designed to detect dark matter with a 35-50 kg NaI(Tl) target and an active background rejection system. The apparatus is made up of three sub-detector systems: (i) the NaI(Tl) crystal detector array, (ii) the linear alkylbenzene (LAB) liquid scintillator veto system, and (iii) the plastic scintillator muon veto system. Together, the liquid scintillator and muon detectors act as an active veto, rejecting events from intrinsic and extrinsic background sources. These components are shown in Figs. 4 and are detailed in the following sections.
@@ 4 Crystal modules
The NaI(Tl) crystal detector system is the core of the detector, and significant effort has been put into photomultiplier tube (PMT) characterisation, crystal growth, enclosure design, and insertion system design. The crystals are grown from Merck (formerly Sigma-Aldrich) Astrograde NaI powder, which has potassium contamination below 10 ppb [18]. Upon arrival at SUPL, they are stored underground and assembled in a nitrogen-flushed assembly glove box, housed in a radon- and dust-reduced space.
SABRE collaborates with SICCAS (China) and RMD (USA) to produce ultra-pure NaI(Tl) crystals. RMD has produced several prototypes that were tested at LNGS (Laboratori Nazionali del Gran Sasso) for optical and radiopurity properties. SICCAS is currently growing additional test crystals. Low levels of natural radiation within the NaI powder and other detector materials constitute a key background component. These can produce energy depositions in the region of interest (1-6 keV) that obscure the DM signal. The isotopes that form significant background near the region of interest are 40K and 210Pb, alongside others depicted in Fig. 3. Note that the 87Rb contribution is from an upper limit measurement and is expected to be significantly lower in practice.
Each crystal is coupled to two 76 mm-diameter Hamamatsu R11065 PMTs, and is enclosed in an oxygen-free, high-thermal-conductivity (OFHC) copper enclosure. These crystal enclosures are then submerged in the SABRE South vessel, suspended by copper conduits that house the cables from the PMTs inside the vessel. They have been fabricated at the Australian National University and have been designed to be leak-proof when immersed in LAB for the duration of the experiment. To ensure a moisture-, oxygen-, and radon-free volume, the enclosures will also be flushed with high-purity nitrogen, the handling and control of which will be managed by a bespoke gas handling system (see Sec. 4.5).
@@
NaI(Tl) crystals 7 $\times$ 5-7 kg - length $\sim$15-20 cm, dia. $\sim$10 cm
@@ 4.1 Enclosure design
SABRE South supports seven crystal detector modules. The enclosures are designed to ensure the aforementioned contaminant-free environment during handling, assembly, and insertion of the NaI(Tl) detector modules. A render of a crystal enclosure, with the associated conduit/support pipe, is shown in Fig. 5. Each enclosure is suspended by an enclosure support pipe that is connected to the top endcap in the top flange of the veto vessel. It carries the PMT signal and high voltage (HV) connections, gas lines, and other sensor cables, while keeping them isolated from the liquid scintillator (LS).
@@
[…]
The enclosures are made entirely of OFHC copper (C0100) and PTFE. The OFHC copper components are extremely radiopure while being commercially available. The outer shell of the crystal enclosure has an outer diameter of 143 mm, a maximum wall thickness of 3 mm and a length of 670 mm. The hermetic seals of the enclosure are achieved using ‘V-grooves’ that compress O-rings with a 45 chamber in the endcap plates against the copper shell.
@@
Shown and fully labelled in Fig. 6, the internal crystal sub-assembly is designed to mechanically support the NaI(Tl) crystal and the coupled PMTs within the copper shell, ensuring that coupling is maintained during experimental run time and through the later stages of assembly. The crystal is held with two PTFE holders that sit above and below the crystal with a cutout for the PMTs. These are connected with three OFHC-copper compression rods that provide mechanical support for the crystal-PMT coupling. Each PMT is held in place with a copper and PTFE holder, which is connected to the compression rods. This design provides an evenly distributed pressure on the PMTs and the optical grease, which ensures good optical coupling to the NaI(Tl) crystals without causing excessive stress on the PMTs. The compression rods are 7 mm in diameter with partial threading to attach the PMT holders, to allow for variations in crystal height.
The crystal sub-assembly is connected to the enclosure endcaps with three copper columns that, when assembled, provide the mechanical support for the sub-assembly, shown in Fig. 5. As the crystal growth process may lead to crystals of differing dimensions, the crystal enclosure has been designed to accommodate variations both in crystal height and diameter. Consequently, there are components that are either independent of or dependent on the size of the crystal. The main components dependent on size are the compression rods and copper support columns. These are made to the crystal dimensions and are designed to align the crystal and veto vessel midplanes. Crystals with a height less than the limit will have more clearance at each end. The remaining components are independent of crystal size and allow for modular manufacturing and assembly prior to the crystal growth being finalised.
closure has been constructed with this design,edct assembly tests to confirm the suitability of the design. In addition, a top endcap was manufactured to conduct leak testing using a helium leak checker, which will be repeated for each manufactured enclosure. The crystal enclosure components independent of crystal size have since been manufactured.

### v3

@@ 1 Introduction
Summarised in this document is the Technical Design Report for the SABRE (Sodium Iodide with Active Background REjection) South experiment. SABRE South is a NaI(Tl) dark matter direct detection experiment, to be situated underground in the Stawell Underground Physics Laboratory in the Southern Hemisphere, and aims to test the results from the DAMA/LIBRA experiment. SABRE South’s three key sub-detectors are described, alongside other major components of the experiment design such as data acquisition, and shielding. The expected sensitivity of the detector to a DAMA/LIBRA-like signal is also discussed.
@@ 1.1 Scientific goals
A wide range of astrophysical evidence supports the existence of non-luminous matter known as dark matter (DM) [1, 2, 3, 4, 5]. Based on these observations, dark matter must have the following properties: it interacts gravitationally, but is electrically neutral and does not interact via the strong force; it is largely non-relativistic (“cold”); it is around five times more abundant than baryonic matter; it has a freeze-out time that allows for the observed structural evolution of the universe; and, it is stable on cosmological timescales. Any observable signal from direct detection should have a distinct modulation signature due to the Earth’s revolution around the Sun as it moves through the galactic dark matter halo [2]. This modulation is unrelated to the Earth’s seasons and is identical in both hemispheres. To be compatible with DM, the modulation signal should satisfy the following conditions [6, 2]: the modulation should have a period of one year; the modulation’s phase should produce a peak at the start of June; for GeV-scale dark matter particle(s), the modulation should be present in a low energy range (keV-scale); the dark matter particle interacts rarely with matter, expecting only single hits with the target material; the modulation amplitude is not larger than $\sim$30of the total dark matter signal (the exact amount varies depending on the particle interaction model).
@@
The SABRE South detector is also capable of testing other models of dark matter and their interactions, as well as detecting supernova neutrinos out to the galactic centre, as secondary goals. Preliminary sensitivity studies have been performed for the detection of spin-dependent WIMPs via the Migdal effect [12], the detection of bosonic super WIMPs [13], and for the use of the SABRE South detector in a supernovae early warning system (SNEWS) [14]. Utilising the SABRE South liquid scintillator veto, we would expect to observe approximately 6 events from a typical core-collapse supernova event at 10 kpc, assuming a 3-fold coincidence requirement and the nominal number of PMTs. While the number of detection events would be smaller than other detectors in the SNEWS system, an additional experiment in the Southern Hemisphere, would be seen as valuable. Publications discussing the SABRE South detector’s sensitivity to various dark matter models, and to supernovae neutrinos, are expected in the future.
@@ 1.2 Dark Matter searches in NaI(Tl) detectors
DAMA/LIBRA, like its predecessor DAMA/NaI, aims to observe the model-independent modulation of a dark matter interaction rate. The most recent iteration, DAMA/LIBRA-phase2 used 250 kg of NaI(Tl) over 8 years of data collection. Combining this with previous experiments, DAMA/LIBRA has a total exposure period of 2.86 tonne years. The interaction rate itself, integrated over a given energy window $k$, can be given by
@@
where $R_{0,k}$ is the constant rate, $S_{m,k}$ is the amplitude of the modulating component of the rate, $\omega$ is the angular frequency, and $t_{0}$ is the phase [15]. Figure 1 depicts how the modulation is induced by the Earth’s motion through the WIMP wind.
Figure 1: A diagrammatic depiction of the Earth’s motion through the WIMP galactic halo (the WIMP wind effect), and the times of year at which the rate is at a maximum and at a minimum.
@@
The DAMA collaboration has consistently observed a modulating signal over their background rate of 0.8 cpd/kg/keV [15] (counts per day/kg/keV) throughout the DAMA/NaI, DAMA/LIBRA-phase1 and DAMA/LIBRA-phase2 experiments in the 2–6 keV region, which has a modulation amplitude of $0.01014\pm 0.00074$ cpd/kg/keV with a phase and period of 142.4 $\pm$ 4.2 days and 0.99834 $\pm$ 0.00067 years respectively [16]. In the 1–6 keV recoil energy range (reported for DAMA/LIBRA-phase2 only) this modulation occurs with an amplitude of $0.01058\pm 0.00090$ cpd/kg/keV, period of $0.99882\pm 0.00065$ years and phase of $144.5\pm 5.1$ days. Recent preliminary results indicate that this modulation persists below the 1 keV threshold [8].
@@
The signal observed by DAMA/LIBRA is in direct tension with a wide variety of direct-detection dark matter experiments using different target materials [17]. Recently, there have also been studies suggesting that DAMA/LIBRA’s background averaging technique (if performed over a time-dependent background) could have induced a modulating result [18]. However, after reproducing a best estimate of their background, we find that: (i) the reproduced background is too low for a modulation to be induced on the same scale as their reported signal, and (ii) the resulting induced modulation is opposite in phase to the modulation reported by DAMA/LIBRA [19]. Therefore, in recent years, other NaI(Tl)-based experiments have been developed to test this modulation using the same target as DAMA, such as COSINE [20] and ANAIS [21]. We list the observed modulation from each of the operating experiments in Table 1. Results in the 1–6 keV region ANAIS are known to be impacted by higher than expected background contributions that are unexplained [22, 21]; a potential cause of this is photomultiplier noise [23]. Additionally, the signal efficiencies of the event selection criteria drop steeply as 1 keV is approached. For example, between 1–2 keV, the ANAIS detection efficiency drops to as low as 20%, with the degree of background rejection not reported [21]. For COSINE, the detection efficiency is around 40% at 1 keV, with a sharp efficiency turn-on effect over the energy interval of 1–2 keV [24], which may lead to non-negligible systematic uncertainties. As a result, a comparison of results in the 2–6 keV region is likely more robust. To definitively test DAMA/LIBRA, SABRE must be sensitive to a modulation of 0.01 cpd/kg/keV, which requires a total background of no more than 1 cpd/kg/keV, and ideally better than the DAMA value of 0.8 cpd/kg/keV [15].
@@
Table 1: Annual modulation results reported by DAMA [16], COSINE [20, 22], and ANAIS [25] (ANAIS also reports updated results from their 3-year data [21]). We define dru (differential rate units) as cpd/kg/keV.
Setup Mass (kg) Bkg. (dru) $S_{m}$ (1–6 keV) (dru) $S_{m}$ (2–6 keV) (dru)
DAMA/LIBRA-phase2 250 0.8 0.0105 $\pm$ 0.0009 0.0093 $\pm$ 0.0009
COSINE 61.3 2.7 0.0017 $\pm$ 0.0029 0.0051 $\pm$ 0.0047
@@ 2 The Stawell Underground Physics Laboratory
Figure 2: Left: location of SUPL. Right: muon flux at a number of underground laboratories as a function of water equivalent depth. The point for SUPL is based on simulated results using GEANT4 [26] and MUTE [27]. The points denote flux measurements in each laboratory. The dashed lines are fits to the data points according to the method laid out in Ref. [28]. The black line and points represent labs with a flat overburden, and the blue line and points represent labs with mountainous overburden (with the blue region denoting the error on the fit). Not included in the figure is the data for Y2L: publicly available data indicates that Y2L has depth of about 2.4 km of water equivalent with mountainous overburden [29].
The construction of SUPL was completed in 2023, with the first access for SABRE South installation activities in early 2024. The assembly and commissioning of the main detector components in SUPL is currently scheduled for 2025, after which data collection will begin. SUPL has helical drive access through mining tunnels and no vertical access shafts, which limits the size and weight of detector components. The Stawell Gold Mine remains an active gold mine; however, SUPL is located in a disused portion of the mine.
@@ 3 The SABRE South concept
To definitively test the DAMA/LIBRA anomaly, it is essential to use the same target material and achieve an ultra-low background to detect the reported modulation amplitude of 0.01 cpd/kg/keV. Since the primary contributor to the total background is the radiogenic background in the crystals, the SABRE South experiment aims to achieve background levels comparable to those of DAMA/LIBRA by ensuring high NaI(Tl) radiopurity. Additionally, by employing a background veto system—comprising the liquid scintillator detector and the muon detector—external background and residual radioimpurities can be tagged and removed [30].
The test crystals grown for the SABRE collaboration are very competitive with other NaI(Tl) crystals in terms of radiopurity [31, 32, 33, 34]. SABRE South’s location in the Southern Hemisphere also offers a unique advantage by providing data that can disentangle seasonal effects.
@@ 3.1 Radiopurity of NaI(Tl) detectors
SABRE South and SABRE North are designed such that radiogenic background intrinsic to the crystals dominate the total background, as demonstrated through simulations [11, 35] and indicated in Refs. [32, 31]. This design highlights that radiopure crystals are the key factor in achieving an experiment capable of reproducing the DAMA/LIBRA sensitivity. The purity levels of the DAMA, ANAIS, COSINE, and SABRE crystals are listed in Table 2, with the SABRE South crystal radiogenic background spectrum (Fig. 4) taken from the measurements in Ref. [34], which utilise the SABRE Proof-of-Principal crystal. The veto efficiency is applied. The SABRE crystals have a 40K contamination better than that of DAMA/LIBRA and much lower than than the other existing NaI(Tl) experiments (COSINE and ANAIS). Another important contaminant is 210Pb that can accidentally enter the crystal during powder production and handling, or in the growth procedure, due to exposure to  [^2] Rn [34, 32, 36]. The background spectra in Fig. 4, including the 210Pb spectrum, are based on measurements reported in Ref. [34]. The 210Pb background spectrum derives from concentrations measured after the crystal was chemically polished with semiconductor-grade ethanol/isopropyl alcohol.
As shown in Ref. [34], SABRE crystals have the lowest NaI(Tl) background worldwide, enabling SABRE to achieve the 0.72 cpd/kg/keV total background reported in Ref. [11]. Further details on NaI(Tl) background modelling for SABRE South can be found in Ref. [11], and crystal characterisation is discussed in Ref. [34].
SABRE South is considering two different final scenarios for crystal growth: one with a total mass of approximately 50 kg (7$\times$7 kg crystals grown using the Bridgman method by the Shanghai Institute of Ceramics, Chinese Academy of Sciences (SICCAS)), and another with a total mass of 35 kg (7$\times$5 kg crystals grown using zone refining followed by the Bridgman method by Radiation Monitoring Devices (RMD), Boston). References to a total active mass of 35-50 kg reflect these different scenarios. In both cases the Bridgman method is chosen, as opposed to the Kyropoulos technique, as the molten raw material can be completely sealed inside the crucible during crystal growth, reducing the risk of contamination.
@@
SABRE (NaI-033) [33, 38] $4.3$ $0.4$ $0.2$ $0.5$
Figure 4: The crystal background rate from intrinsic contamination at SABRE South for 50 kg of NaI(Tl) crystals [11]. The units keVee refer to electron equivalent energy. These spectra are based on measurements of the SABRE Proof-of-Principle crystal [34]. Note: The y-axis on this does not extend down to zero, and the spectra for 129I and 232Th continue below the lower y-axis limit. Details on their total background rate below this limit can be found in Ref. [11].
@@ 3.2 SABRE South detector
The SABRE South experiment is designed to detect dark matter with a 35-50 kg NaI(Tl) target and an active background rejection system. The apparatus is made up of three sub-detector systems: (i) the NaI(Tl) crystal detector array, (ii) the linear alkylbenzene (LAB) liquid scintillator veto system, and (iii) the plastic scintillator muon veto system. Together, the liquid scintillator and muon detectors act as an active veto, rejecting events from intrinsic and extrinsic background sources. These components are shown in Fig. 5 and are detailed in the following sections.
@@ 4 Crystal modules
The NaI(Tl) crystal detector system is the core of the SABRE South detector, and significant effort has been put into photomultiplier tube (PMT) characterisation, crystal growth, enclosure design, and insertion system design. The crystals are grown from Merck (formerly Sigma-Aldrich) Astrograde NaI powder, which has potassium contamination below 10 ppb [39]. Upon arrival at SUPL, they are stored underground and assembled in a nitrogen-flushed assembly glove box, housed in a radon- and dust-reduced space.
SABRE collaborates with SICCAS (China) and RMD (USA) to produce ultra-pure NaI(Tl) crystals. RMD has produced several prototypes that were tested at LNGS (Laboratori Nazionali del Gran Sasso) for optical and radiopurity properties. SICCAS is currently growing additional test crystals. Low levels of natural radiation within the NaI powder. These can produce energy depositions in the region of interest (1–6 keV) that obscure the DM signal. The isotopes that form a significant background near the region of interest are 40K and 210Pb, alongside others depicted in Fig. 4. Note that the 87Rb contribution is from an upper limit measurement and is expected to be significantly lower in practice. We plan to measure the quenching factors (QFs) of the crystals, using the tips and tails of each crystal ingot, as the QFs may vary from crystal to crystal. SABRE South has performed a QF measurement on a test NaI(Tl) crystal in Ref. [40], which details the procedure that we plan to follow for later measurements.
Each crystal is coupled with optical grease to two 76 mm-diameter Hamamatsu R11065 PMTs, and is enclosed in an oxygen-free, high-thermal-conductivity (OFHC) copper enclosure. The sides of the crystals are wrapped in PTFE reflector to ensure optimal light collection. The PTFE can contain 210Pb contaminants. The contribution of this radioimpurity to the background depends on whether it is on the surface or in the bulk of the material. Deposits close to or on the surface will lead to higher background levels [38]. Improved simulations of light collection efficiency are being developed and will be compared with upcoming test measurements at SUPL. These crystal enclosures are then submerged in the SABRE South vessel, suspended by copper conduits that house the cables from the PMTs inside the vessel. The crystals are placed vertically rather than horizontally to ensure that they do not slide against the PMTs under gravity, and for practicality, as a smaller opening in the vessel is required for insertion. They have been fabricated at the Australian National University and have been designed to be leak-proof when immersed in LAB for the duration of the experiment. To ensure a moisture-, oxygen-, and radon-free volume, the enclosures are flushed at a low flow rate with high-purity nitrogen, the handling and control of which is managed by a bespoke gas handling system (see Sec. 4.5).
@@
NaI(Tl) crystals 7 $\times$ 5–7 kg - length $\sim$15–20 cm, dia. $\sim$10 cm
@@ 4.1 Enclosure design
SABRE South supports seven crystal detector modules. The enclosures are designed to ensure the aforementioned contaminant-free environment during handling, assembly, and insertion of the NaI(Tl) detector modules. A render of a crystal enclosure, with the associated conduit/support pipe, is shown in Fig. 6. Each enclosure is suspended by an enclosure support pipe that is connected to the top endcap in the top flange of the veto vessel. It carries the PMT signal and high voltage (HV) connections, gas lines, and other sensor cables, while keeping them isolated from the liquid scintillator.
@@
The enclosures must meet a strict set of requirements. The leak rate across the sealing surfaces can not exceed $1\times 10^{-6}$ mbar L/s. They are to be immersed in the LAB liquid scintillator for at least 5 years without degradation of the liquid scintillator or ingress of the scintillator into the enclosure. Long-term material compatibility studies performed by SABRE South have confirmed that all materials in direct contact with the liquid scintillator are compatible and are not expected to break down. The enclosures must provide a minimal contribution to the experimental background. The material between the crystals and the liquid scintillator must be minimised to achieve optimal veto efficiency. The dimensions of the enclosures must comply with the restrictions imposed by the crystal insertion system and the veto vessel top flange. The crystal dimensions are expected to vary slightly in size, so the enclosures must be adaptable enough to accommodate a range of sizes. The design allows a height of up to 250 mm, and a diameter limit of approximately 110 mm, corresponding to an 8.72 kg crystal. Finally, the assembly of the modules is performed inside a glovebox, so the design must accommodate assembly with a restricted range of motion.
The enclosures are made entirely of OFHC copper (C0100) and PTFE. The OFHC copper components are extremely radiopure while being commercially available, and were sourced from Thyssenkrupp in the United States. The copper and PTFE components are machined at the Australian National University (ANU). The copper surfaces are chemically etched to reduce 210Pb plate-out and other radioactive impurities, and the PTFE is vacuum baked to reduce out-gassing. The outer shell of the crystal enclosure has an outer diameter of 143 mm, a maximum wall thickness of 3 mm and a length of 670 mm. The hermetic seals of the enclosure are achieved using ‘V-grooves’ that compress O-rings with a 45 chamber in the endcap plates against the copper shell.
@@
Shown and fully labelled in Fig. 7, the internal crystal sub-assembly is designed to mechanically support the NaI(Tl) crystal and the coupled PMTs within the copper shell, ensuring that coupling is maintained during experimental run time and through the later stages of assembly. The crystal is held with two PTFE holders that sit above and below the crystal with a cutout for the PMTs. These are connected with three OFHC-copper compression rods that provide mechanical support for the crystal-PMT coupling. Each PMT is held in place with a copper and PTFE holder and directly coupled to the crystal with optical grease, which is connected to the compression rods. This design provides an evenly distributed pressure on the PMTs and the optical grease, which ensures good optical coupling to the NaI(Tl) crystals without causing excessive stress on the PMTs. The compression rods are 7 mm in diameter with partial threading to attach the PMT holders, to allow for variations in crystal height.
The crystal sub-assembly is connected to the enclosure endcaps with three copper columns that, when assembled, provide the mechanical support for the sub-assembly, shown in Fig. 6. As the crystal growth process may lead to crystals of differing dimensions, the crystal enclosure has been designed to accommodate variations both in crystal height and diameter. Consequently, there are components that are either independent of or dependent on the size of the crystal. The main components dependent on size are the compression rods, copper support columns, and the PTFE crystal holders. These are made to the crystal dimensions and are designed to align the crystal and veto vessel midplanes. Crystals with a height less than the limit will have more clearance at each end. The remaining components are independent of crystal size and allow for modular manufacturing and assembly prior to the crystal growth being finalised.
A prototype crystal enclosure has been constructed with this design, and was used to conduct assembly tests to confirm the suitability of the design. In addition, a top endcap was manufactured to conduct leak testing using a helium leak checker, which will be repeated for each manufactured enclosure. The crystal enclosure components independent of crystal size have since been manufactured.